# DREAMS: A python framework to train deep learning models with model card reporting for medical and health applications

### v1

@@ 5 Impact
6 Conclusions and future perspectives
This paper introduces DREAMS, a Python library designed to simplify the creation of comprehensive model cards for deep learning models.
In fact, DREAMS offers efficient generation of detailed model cards with customizable content through YAML configuration, covering project overviews, dataset descriptions, results with plots, uncertainty estimations, and references. It also supports various plot types, enhancing clarity and comprehension.
As such, DREAMS serves as a valuable tool for promoting transparency and ethical use of AI across various domains, facilitating informed decision-making, and fostering trust in AI systems. Whether in healthcare, finance, intelligent transportation systems, or any other domain where AI is utilized, DREAMS enables the documentation of model performance, usage, and limitations through comprehensive model cards. This documentation not only enhances understanding but also promotes accountability and ensures that AI systems are deployed responsibly and ethically.

### v2

@@ 6 Impacts
7 Discussions and Future Work
This work introduces DREAMS, a Python-based model card framework for EEG-based deep learning models, designed to enhance transparency, interpretability, and reproducibility. A key advantage of this framework is the integration of automated model documentation, reducing manual effort while ensuring consistency across model iterations. The incorporation of confidence intervals in model evaluation improves interpretability, enabling researchers to quantify prediction reliability and assess model uncertainty. Furthermore, dataset characteristics, model´s hyperparameters, preprocessing details, and performance analysis can be systematically documented, ensuring comprehensive and standardized reporting across studies.
Beyond improving reproducibility, DREAMS also addresses critical ethical concerns in EEG-based AI by promoting responsible model documentation. Ethical challenges in deep learning, particularly in EEG-based healthcare and neuroscience applications, include bias in training data, lack of transparency in decision-making, and potential risks in clinical deployment. By explicitly documenting dataset demographics, potential biases, and model limitations, the model card enables researchers to assess fairness and mitigate algorithmic bias. Additionally, DREAMS allows for the inclusion of ethical guidelines, ensuring that researchers disclose data handling practices, consent considerations, and limitations of model generalizability, fostering greater accountability in AI-driven EEG applications.
The customizability of DREAMS ensures that model cards can be extended based on specific use cases, incorporating additional sections such as deployment considerations, ethical guidelines, and regulatory compliance requirements. This flexibility allows researchers to tailor documentation for clinical validation, healthcare deployment, or domain-specific performance evaluation, supporting ethical AI development in high-stakes applications.
While model documentation is a significant step toward improving reproducibility, broader challenges remain in EEG-based deep learning. The variability of EEG signals across individuals introduces inherent uncertainty, making generalization difficult. This necessitates future work on subject-adaptive models and personalized EEG feature extraction. Additionally, the lack of standardized reporting practices in EEG-based deep learning further underscores the need for structured model cards as part of a transparent and ethical AI workflow. By systematically addressing uncertainty, bias, and documentation gaps, DREAMS provides a foundational framework for responsible EEG-based AI research and deployment.
Future research should explore the integration of explainability techniques, such as saliency maps and attention mechanisms, to provide deeper insights into EEG feature learning. Furthermore, ensuring fairness and bias detection in EEG models, particularly across different demographics, remains an important direction for future study.
In conclusion, this work provides a structured and automated model card framework designed to improve reproducibility, interpretability, and uncertainty-aware evaluation in EEG-based deep learning. We show the integration of DREAMS into deep learning workflow to enhance model transparency by systematically reporting performance metrics such as accuracy, precision, recall, and F1-score with 95% confidence intervals. DREAMS ensures automated documentation with version tracking, reducing manual effort and improving reproducibility across model iterations. By addressing the lack of standardized model documentation, our framework ensures that performance variability, dataset characteristics, and model reliability are systematically reported. It can be extended to various EEG classification tasks, clinical research, and neuroscience applications. Through the introduction of this framework, we aim to establish a standard for transparent and reproducible model reporting, ultimately contributing to responsible AI adoption in EEG research and clinical settings.